OpenStudy (anonymous):

what is the fourth term of (x-2)^8

mathslover (mathslover):

Do you know the formula for getting a term in binomial expansion?

OpenStudy (anonymous):

use pascal's triangle.

mathslover (mathslover):

No need to do that mack.

mathslover (mathslover):

Here is the formula to find the (n+1)th term of expression : \((x+y)^a\) \(\large T_{n+1} = ^a C _ n x^{n-a} y^a \) \ Here we are given with the expression : \((x-2)^8\) , that is a = 8 \(\large T_{4} = T_{3+1} = ^8 C _ 3 x^{8-3} (-2)^3 \) [ By : \(T_n \) I meant, nth term. ]

mathslover (mathslover):

Solving it further, you will get the 4th term of the expansion : \((x-2)^8\)